OpenStudy (anonymous):

Please help Let f(x)=2x^2+x-3 and g(x)=x-1. Perform the indicated operation, then find the domain. (F-g)(x) A. X^2-4; domain: positive real numbers B. 2x^2-4; domain: all real numbers C. X^2-4; domain: all real numbers D. 2x^2-2; domain: all real numbers

OpenStudy (anonymous):

@mathmath333 @mathmate

OpenStudy (mathmate):

Can you do the (F-g)(x) part?

OpenStudy (anonymous):

No:/, that's why I posted the question on here @mathmate

OpenStudy (mathmate):

(f-g)(x) = f(x)-g(x) so you would subtract the polynomial g(x) from f(x).

OpenStudy (mathmate):

In case of need, you can use the following link to help you, it is easy to read and full of examples: http://www.purplemath.com/modules/polyadd2.htm

OpenStudy (mathmate):

gtg. two more things that will help you: 1. Sum and difference of polynomials are polynomials. 2. All polynomials have a domain of real numbers, i.e. \((-\infty,\infty)\)